OpenStudy (anonymous):

Find an explicit rule for the nth term of the sequence. 2, -8, 32, -128, ...

OpenStudy (jdoe0001):

so, what do you think is the common ratio?

OpenStudy (anonymous):

*-4

OpenStudy (jdoe0001):

hmm, well, I guess that'd be the actual query.. so in this geometric sequence is just a multiplier btw \(\begin{array}{ccccc} 1^{st}&2^{nd}&3^{rd}&4^{th}\\ \hline\\ 2,& -8,& 32,& -128,& ....\\ \hline\\ 4^{1-1}&4^{2-1}&4^{3-1}&4^{4-1} \end{array}\)

OpenStudy (jdoe0001):

woops, should be ... -4 rather :(

OpenStudy (jdoe0001):

\( \begin{array}{ccccc} 1^{st}&2^{nd}&3^{rd}&4^{th}\\ \hline\\ 2,& -8,& 32,& -128,& ....\\ \hline\\ -4^{1-1}&-4^{2-1}&-4^{3-1}&4^{-4-1} \end{array}\) so as you can see, the rule will be \(\bf \large -4^{n-1}\) , n= term ordinal position

OpenStudy (jdoe0001):

shoot I stil have a typo... ohh, well anyhow hehe

OpenStudy (anonymous):

so it would be 2*-4^n-1

OpenStudy (anonymous):

lol that's okay xD I am understanding

OpenStudy (jdoe0001):

hmm .... yeah... I sorta skipped the a1

OpenStudy (anonymous):

thanks :)

OpenStudy (jdoe0001):

yw